OpenStudy (anonymous):

Find the 24th term in the sequence for which a=-27 and d=3? how do you plug that in?

OpenStudy (anonymous):

so 24th term means N = 24 , a = -27 and d = 3. Nth term is given By Tn = a + (n-1)d,

OpenStudy (anonymous):

so An=- 27 + (24-1)3?

OpenStudy (anonymous):

A24 = -27 + (24-1)3 yes thats correct

OpenStudy (anonymous):

Thanks just making sure that was correct